OpenStudy (anonymous):

I JUST NEED THE ANSWER!!!!! Find the product of the solutions of this equation. 15X^2+23X-41=0

OpenStudy (anonymous):

Use the quadratic formula to find the solutions. Do you know the quadratic formula?

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then product of solutions will be c/a.

OpenStudy (anonymous):

Y a but it will take to long I need the answer fast...

OpenStudy (anonymous):

and its hard

OpenStudy (anonymous):

It's not very difficult. Do you have access to a calculator?

hartnn (hartnn):

@BlizzardofFire could u find a,b,c ?

OpenStudy (anonymous):

@BlizzardofFire i gave you the answer...

OpenStudy (anonymous):

ya... but i have to answer 25 more math questions in 30 min so i dont have alot of time

OpenStudy (anonymous):

@Turner, you gave him the solutions to the equation. Not the answer. The answer is the product of the two solutions.

hartnn (hartnn):

a=15,c=-41 so product of solution = c/a = -41/15.

OpenStudy (anonymous):

@Hartnn, also incorrect.

OpenStudy (anonymous):

thanks hartnn

hartnn (hartnn):

what ? @teic85 are u sure i am incorrect ?

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

First of all, c/a is a quotient, not a product

hartnn (hartnn):

then check your solution, what did u get ?

OpenStudy (anonymous):

I haven't solved it. I'm only here to provide a path to the answer. Anything else is academically unethical. Plus, no one learns anything from just being given the answer.

OpenStudy (anonymous):

Can you explain to me how you arrived at your solution?

OpenStudy (anonymous):

it was one of the choices i have a new question up

OpenStudy (anonymous):

Are you taking a test?